# Semiparametric Copula Estimation for Spatially Correlated Multivariate Mixed Outcomes: Analyzing Visual Sightings of Fin Whales from Line Transect Survey

### v2

@@ Introduction
For the simulated dataset, we apply the proposed and Hoff (2007) methods, denoted as spBGC and BGC, respectively. In doing so, we use 2000 draws for the posterior computation after discarding the first 1000 draws as burn-in, and we set $IW(p+2,(p+2)\bm{I})$ prior for $\bm{V}$ as a prior for $\bm{V}$. We compute posterior medians $\{\hat{\bm{R}}_{jj^{\prime}}\}_{j<j^{\prime}}$ as point estimates of $\bm{R}_{jj^{\prime}}$ for $j<j^{\prime}$ and evaluate their performance by the mean squared error (MSE), defined as $q^{-1}\sum_{j<j^{\prime}}(\hat{\bm{R}}_{jj^{\prime}}-\bm{R}_{jj^{\prime}})^{2}$ where $q=p(p-1)/2$. We also compute 95% credible intervals and calculate coverage probabilities (CP) and average lengths (AL), defined as $q^{-1}\sum_{j<j^{\prime}}I(\bm{R}_{jj^{\prime}}\in{\rm CI}_{jj^{\prime}})$ and $q^{-1}\sum_{j<j^{\prime}}|{\rm CI}_{jj^{\prime}}|$, respectively, where $I(\cdot)$ is the indicator function and ${\rm CI}_{jj^{\prime}}$ is the 95% credible interval of $\bm{R}_{jj^{\prime}}$. These values are averaged over 300 replications of the simulated datasets.
@@
Furthermore, we determine the advantages of using NNGP in the spBGC. Hereafter, we denote spBGCNNGP to clarify its use. As described below, although spBGCNNGP can considerably reduce the computation time compared to spBGC, even for a sample size of 500, its estimation accuracy remains almost the same. Because spBGC requires considerable time per simulation when the sample size is 1000, we evaluate the performance of spBGC with only 10 iterations in this setting. This simulation study was conducted on an Apple Mac Studio with an Apple M1 Ultra chip, using programming language R (R Core Team, 2024).
@@
The simulation results presented in Figure 2 and Tables 14 clearly demonstrate the superior performance of our proposed methods, spBGC, and its computationally efficient version, spBGCNNGP, compared with the existing BGC method. Figure 2(a) and Table 1 show that both spBGC and spBGCNNGP achieve significantly lower MSEs than BGC across all scenarios. This difference becomes more pronounced as spatial correlation ($\phi$) increases, highlighting the ability of the proposed method to effectively account for spatial dependence. The CPs displayed in Figure 2(b) and Table 2 reveal that spBGC and spBGCNNGP consistently maintain CPs close to the nominal 95% level. By contrast, BGC coverage deteriorates substantially as the spatial correlation strengthens, often falling well below the desired level. This underscores the reliability of the proposed method for providing accurate credible intervals. The ALs of the credible intervals, as shown in Figure 2(c) and Table 3, indicate that spBGC and spBGCNNGP achieve high coverage probabilities, while maintaining reasonably narrow interval widths. Balancing coverage and precision is crucial for practical applications. Most notably, Table 4 highlights the computational efficiency of spBGCNNGP. As the sample size increases, the time savings offered by spBGCNNGP become increasingly significant. For instance, for $n=1000$, spBGCNNGP completes the computation in approximately one-sixth of the time required by the spBGC, while maintaining a comparable estimation accuracy.
@@
These results collectively demonstrate that the proposed methods, particularly the spBGCNNGP, offer a powerful combination of estimation accuracy and computational efficiency for analyzing spatially correlated multivariate mixed outcomes. The ability of spBGCNNGP to significantly reduce computation time without sacrificing estimation quality is particularly valuable for large-scale spatial datasets, which are becoming increasingly common in various research areas. By employing the spBGCNNGP, researchers and practitioners can perform high-precision estimations of spatial dependence structures on datasets that were previously computationally infeasible. This opens new possibilities for in-depth spatial analyses across a wide range of applications, from environmental science to epidemiology. In conclusion, the spBGCNNGP is a robust and efficient tool for analyzing spatially correlated multivariate mixed outcomes, offering an optimal balance between accuracy and computational efficiency. Its ability to handle large-scale spatial datasets with high precision renders it an invaluable asset for researchers studying complex spatial datasets.
@@
Table 1: Comparisons of the logarithm of the MSEs for the spBGC, spBGCNNGP, and BGC under varying sample sizes ($n\in\{50,500,1000\}$) and spatial range parameters ($\phi\in\{0.05,0.25,0.5\}$). The values represent average log(MSE)s from 300 calculations, with standard errors between parentheses.
@@
spBGC $-6.596$ (0.063) $-6.485$ (0.072) $-6.053$ (0.156)
spBGCNNGP $-6.673$ (0.015) $-6.365$ (0.018) $-6.032$ (0.024)
BGC $-5.465$ (0.016) $-3.324$ (0.017) $-2.834$ (0.019)
Table 2: Comparison of CPs for the spBGC, spBGCNNGP, and BGC across varying sample sizes ($n\in\{50,500,1000\}$) and spatial range parameters ($\phi\in\{0.05,0.25,0.5\}$). The values represent average CPs from 300 calculations, with standard errors between parentheses.
@@
spBGC 0.931 (0.011) 0.956 (0.007) 0.936 (0.012)
spBGCNNGP 0.941 (0.002) 0.945 (0.002) 0.938 (0.002)
BGC 0.669 (0.005) 0.256 (0.005) 0.205 (0.005)
Table 3: Comparison of ALs for the spBGC, spBGCNNGP, and BGC across varying sample sizes ($n\in\{50,500,1000\}$) and spatial range parameters ($\phi\in\{0.05,0.25,0.5\}$). The values represent average ALs from 300 calculations, with standard errors between parentheses.
@@
Table 4: Comparison of execution times for the spBGC and spBGCNNGP across varying sample sizes ($n\in\{50,500,1000\}$) and spatial range parameters ($\phi\in\{0.05,0.25,0.5\}$). The values represent average execution times from 300 calculations, with standard errors between parentheses.
@@
Real data applications
We applied our method to the visual sightings of fin whales from line transect surveys in the Southern Ocean. Visual observations were performed on three of the six vessels participating in the 2019 Area 48 Survey for Antarctic krill (Krafft et al., 2021): R/V Kronprins Haakon (KPH), F/V Cabo de Hornos (CDH), and RRS Discovery (DIS). The details of the observation methods and protocols for fin whales can be found in Biuw et al. (2024). The observed data in this study are the number of fin whales and the acoustic data of krill from vessels operating along the transects. Krill biomass estimates were recalculated based on the acoustic data (details in the Supplementary Material of Krafft et al., 2021). The observation transects were split into 1-nm long segments, resulting in 3,833 segments based on 42 transects, and the number of fin sightings (groups and individuals) and krill biomass were summarized for each segment.
We also included surface temperature (Celsius) and water depth (m) for a 1-nm segment to investigate the association of environmental data with the biological community. Figure 3 shows the spatial plots for krill biomass (Krill), sighting data of fin whales (Whale), surface temperature (SST), depth data (Depth), slope of the depth (Slope), and gradient surface temperature (SST.grd). Brighter/darker dots’ colors for surface temperature and water depth indicate higher/lower temperatures and shallower/deeper water depths, respectively. Water depth data were obtained from the ETOPO 1 bathymetric dataset, available at https://www.ncei.noaa.gov/products/etopo-global-relief-model. Data were extracted for the middle position of each 1-nm segment throughout the survey tracks. Data on SST were obtained from the OISST dataset, available at https://www.ncei.noaa.gov/products/optimum-interpolation-sst, being extracted in the same way as for depth data. The slope of the depth is the maximum rate of change in depth from that cell to its neighbors, calculated using the Slope tool of the Surface toolset in ArcGIS (https://desktop.arcgis.com/en/arcmap/10.3/tools/spatial-analyst-toolbox/an-overview-of-the-surface-tools.htm). The lower the slope value is, the flatter is the terrain, and vice versa. The gradient surface temperature was calculated using the Slope applied to the SST.
(a) Krill
(b) Whale
(c) SST
(d) Depth
(e) Slope
(f) SST.grd
Figure 3: Spatial plots for krill biomass (Krill), sighting data of fin whales (Whale), surface temperature (SST), depth data (Depth), slope of the depth (Slope), and gradient surface temperature (SST.grd).
We are interested in the dependence structure of krill biomass and sighting data of fin whales, and that of each of these outcomes and environmental factors (SST, Depth, Slope, and SST.grd). Therefore, this application focuses on the posterior inference of these dependence structures. We compare the results of the proposed method (spBGC) with those of Hoff (2007)’s method (BGC), which does not consider spatial correlation. Before the comparison, we examine the mixing properties of spBGC. Using an $IW(p+2,(p+2)\bm{I})$ prior, that is, $IW(8,8\bm{I})$ for $\bm{V}$, we conduct MCMC with 25,000 iterations, discarding the first 5,000 as burn-in and retaining every 10th sample from the remaining chain, yielding 2,000 samples for posterior calculations. Figure 4 presents the mixing and autocorrelation results for the correlation coefficients of interest. From these figures, the mixing properties of our spBGC algorithm are quite satisfactory, that is, convergence to stationarity appears and the autocorrelation at lag-20 is close to zero for most elements.
@@
Based on 2,000 posterior draws, we compute the posterior medians and 95% credible intervals of the correlations between the variables, as shown in Figure 5 and Tables 5 and 6. The results reveal that spBGC and BGC yield markedly different estimates for several correlations. In paticular, spBGC intervals suggest that Krill has no strong positive or negative correlation with outcomes other than Whale. By contrast, BGC intervals exclude zero for all Krill- and Whale-related outcomes except Depth. Turning to the structure of conditional independence, we also examined partial correlations. Figure 5 and Tables 5 and 6 also show that these partial correlation estimates differ substantially between spBGC and BGC. Notably, the 95% credible intervals from spBGC exclude only zero for the partial correlation between Krill and Whale, whereas the BGC identifies additional pairs with non-zero partial correlations.
As a related analysis, Solvang et al. (2024) applied standard generalized linear models to the same dataset. The results showed a significant positive association between Krill and Whale. Furthermore, the estimated coefficients on SST were significantly negative, while those on Slope were positive for both Krill and Whale. These results are similar to those obtained for the BGC, possibly because neither BGC nor the generalized linear models account for spatial correlation. In contrast, by using the proposed spBGC that considers spatial correlation, we identified relationships that slightly differ from previous findings. This may provide new interpretations of the relationships between the variables.
@@
SST.grd $-0.0691$ $-0.0318$ $0.0090$
Discussion
@@
The proposed method can be applied and extended in several ways. The first is a Gaussian process formulation for the latent variables, from which we can obtain the posterior predictive distribution of the latent variables in arbitrary spatial locations (including non-sampled locations). Then, using the marginal distribution of each outcome estimated in a parametric or non-parametric manner, we can perform spatial predictions in non-sampled locations. Second, although this study focuses on the latent Gaussian process, the proposed method may be conceptually extended to other copula models, such as elliptical copulas and skew elliptical copulas (Smith, 2021). For example, if one is interested in the tail dependence of a latent process, it may be preferable to consider a $t$-copula model. However, posterior computation of additional parameters using such copula models may not be feasible. In this case, an MCMC sampler based on the Metropolis-Hastings algorithm or more efficient computational algorithms may need to be developed. Third, we extend our method to spatiotemporal data. Multivariate data may contain time and location information. For example, the data used in the application example were observed in the Southern Ocean over different periods. In the analysis of the dependence structure of such data, it may be necessary to consider both spatial and temporal correlations. To this end, we consider a hierarchical Bayesian spatiotemporal model based on extended rank likelihood. However, the development of posterior computation for the model is challenging, thus, a subject to address in future work. Finally, although the proposed methodology is for point-referenced data, it is possible to develop a similar method to estimate the copula for multivariate data observed on a graph.

### v3

@@ 4 Simulation study
For the simulated dataset, we apply the proposed and Hoff (2007) methods, denoted as spBGC and BGC, respectively. In doing so, we use 2000 draws for the posterior computation after discarding the first 1000 draws as burn-in, and we set $IW(p+2,(p+2)\bm{I})$ prior for $\bm{V}$ as a prior for $\bm{V}$. We compute posterior medians $\{\hat{\bm{R}}_{jj^{\prime}}\}_{j<j^{\prime}}$ as point estimates of $\bm{R}_{jj^{\prime}}$ for $j<j^{\prime}$ and evaluate their performance by the mean squared error (MSE), defined as $q^{-1}\sum_{j<j^{\prime}}(\hat{\bm{R}}_{jj^{\prime}}-\bm{R}_{jj^{\prime}})^{2}$ where q=p(p1)/2q=p(p-1)/2. We also compute 95% credible intervals and calculate coverage probabilities (CP) and average lengths (AL), defined as $q^{-1}\sum_{j<j^{\prime}}I(\bm{R}_{jj^{\prime}}\in{\rm CI}_{jj^{\prime}})$ and $q^{-1}\sum_{j<j^{\prime}}|{\rm CI}_{jj^{\prime}}|$, respectively, where $I(\cdot)$ is the indicator function and ${\rm CI}_{jj^{\prime}}$ is the 95% credible interval of $\bm{R}_{jj^{\prime}}$. These values are averaged over 300 replications of the simulated datasets.
@@
Furthermore, we determine the advantages of using NNGP in the spBGC. Hereafter, we denote spBGCNNGP to clarify its use. As described below, although spBGCNNGP can considerably reduce the computation time compared to spBGC, even for a sample size of 500, its estimation accuracy remains almost the same. Because spBGC requires considerable time per simulation when the sample size is 1000, we evaluate the performance of spBGC with only 10 iterations in this setting. This simulation study is conducted on an Apple Mac Studio with an Apple M1 Ultra chip, using programming language R (R Core Team, 2024).
@@
The simulation results presented in Figure 3 and Tables 25 clearly demonstrate the superior performance of our proposed methods, spBGC, and its computationally efficient version, spBGCNNGP, compared with the existing BGC method. Figure 3(a) and Table 2 show that both spBGC and spBGCNNGP achieve significantly lower MSEs than BGC across all scenarios. This difference becomes more pronounced as spatial correlation ($\phi$) increases, highlighting the ability of the proposed method to effectively account for spatial dependence. The CPs displayed in Figure 3(b) and Table 3 reveal that spBGC and spBGCNNGP consistently maintain CPs close to the nominal 95% level. By contrast, BGC coverage deteriorates substantially as the spatial correlation strengthens, often falling well below the desired level. This underscores the reliability of the proposed method for providing accurate credible intervals. The ALs of the credible intervals, as shown in Figure 3(c) and Table 4, indicate that spBGC and spBGCNNGP achieve high coverage probabilities, while maintaining reasonably narrow interval widths. Balancing coverage and precision is crucial for practical applications. Most notably, Table 5 highlights the computational efficiency of spBGCNNGP. As the sample size increases, the time savings offered by spBGCNNGP become increasingly significant. For instance, for $n=1000$, spBGCNNGP completes the computation in approximately one-ninth of the time required by the spBGC, while maintaining a comparable estimation accuracy.
@@
These results collectively demonstrate that the proposed methods, particularly the spBGCNNGP, offer a powerful combination of estimation accuracy and computational efficiency for analyzing spatially correlated multivariate mixed outcomes. The ability of spBGCNNGP to significantly reduce computation time without sacrificing estimation quality is particularly valuable for large-scale spatial datasets, which are becoming increasingly common in various research areas. By employing the spBGCNNGP, researchers and practitioners can perform high-precision estimations of spatial dependence structures on datasets that were previously computationally infeasible. This opens new possibilities for in-depth spatial analyses across a wide range of applications, from environmental science to epidemiology. In conclusion, the spBGCNNGP is a robust and efficient tool for analyzing spatially correlated multivariate mixed outcomes, offering an optimal balance between accuracy and computational efficiency. Its ability to handle large-scale spatial datasets with high precision renders it an invaluable asset for researchers studying complex spatial datasets. To assess the robustness of our findings to the choice of spatial correlation function and copula model, we also conduct simulations using Matérn 3/2 and Matérn 5/2 correlation functions, as well as the $t$-copula model; see Section A of the Supplementary Material.
@@
Table 2: Comparisons of the logarithm of the MSEs for the spBGC, spBGCNNGP (with $m=n/10$), and BGC under varying sample sizes and spatial range parameters. The values represent average log(MSE)s from 300 calculations (10 for spBGC at $n=1000$), with standard errors in parentheses.
@@
spBGC 6.886-6.886 (0.015) 6.393-6.393 (0.017) 6.238-6.238 (0.026)
spBGCNNGP 6.736-6.736 (0.016) 6.352-6.352 (0.018) 6.015-6.015 (0.027)
BGC 5.406-5.406 (0.011) 3.376-3.376 (0.015) 2.860-2.860 (0.016)
Table 3: Comparisons of the coverage probabilities for the spBGC, spBGCNNGP (with $m=n/10$), and BGC under varying sample sizes and spatial range parameters. The values represent average coverage probabilities from 300 calculations (10 for spBGC at $n=1000$), with standard errors in parentheses.
@@
spBGC $0.972$ (0.002) $0.944$ (0.003) $0.956$ (0.003)
spBGCNNGP $0.950$ (0.002) $0.944$ (0.002) $0.941$ (0.002)
BGC $0.664$ (0.005) $0.250$ (0.006) $0.233$ (0.006)
Table 4: Comparisons of the average credible interval lengths for the spBGC, spBGCNNGP (with $m=n/10$), and BGC under varying sample sizes and spatial range parameters. The values represent average interval lengths from 300 calculations (10 for spBGC at $n=1000$), with standard errors in parentheses.
@@
Table 5: Comparison of execution times for the spBGC and spBGCNNGP under varying sample sizes and spatial range parameters. The values represent average execution times from 300 calculations, with standard errors in parentheses.
@@
5 Application to fin whale sighting data
We apply our method to the data set comprising krill biomass data, visual sighting data of fin whales from line transect surveys in the Southern Ocean, and four environmental variables. The detailed description of the data is provided in Section 1.2. We are interested in the dependence structure of krill biomass and sighting data of fin whales, and that of each of these outcomes and environmental factors (SST, Depth, Slope, and SST.grd). Therefore, this application focuses on the posterior inference of these dependence structures. We compare the results of the proposed method (spBGC) with those of Hoff (2007)’s method (BGC), which does not consider spatial correlation. Before the comparison, we examine the mixing properties of spBGC. Using an $IW(p+2,(p+2)\bm{I})$ prior, that is, $IW(8,8\bm{I})$ for $\bm{V}$, we conduct MCMC with 25,000 iterations, discarding the first 5,000 as burn-in and retaining every 10th sample from the remaining chain, yielding 2,000 samples for posterior calculations. For the spatial correlation function $\rho(\cdot;\phi)$ in (3), we use the exponential correlation function. To assess the sensitivity of our results to this choice, we also conduct analyses using Matérn 3/2 and Matérn 5/2 correlation functions; see Section B of the Supplementary Material.
Figure 4 presents the mixing and autocorrelation results for the correlation coefficients of interest. From these figures, the mixing properties of our spBGC algorithm are quite satisfactory, that is, convergence to stationarity appears and the autocorrelation at lag-20 is close to zero for most elements.
@@ 5 Application to fin whale sighting data
Based on 2,000 posterior draws, we compute the posterior medians and 95% credible intervals of the correlations between the variables, as shown in Figure 5 and Tables 6 and 7. The results reveal that spBGC and BGC yield markedly different estimates for several correlations. In paticular, spBGC intervals suggest that Krill has no strong positive or negative correlation with outcomes other than Whale. By contrast, BGC intervals exclude zero for all Krill- and Whale-related outcomes except Depth. Turning to the structure of conditional independence, we also examine partial correlations. Figure 5 and Tables 6 and 7 also show that these partial correlation estimates differ substantially between spBGC and BGC. Notably, the 95% credible intervals from spBGC exclude only zero for the partial correlation between Krill and Whale, whereas the BGC identifies additional pairs with non-zero partial correlations.
As a related analysis, Solvang et al. (2024) applied standard generalized linear models to the same dataset. The results showed a significant positive association between Krill and Whale. Furthermore, the estimated coefficients on SST were significantly negative, while those on Slope were positive for both Krill and Whale. These results are similar to those obtained for the BGC, possibly because neither BGC nor the generalized linear models account for spatial correlation. In contrast, by using the proposed spBGC that considers spatial correlation, we identify relationships that slightly differ from previous findings, with the exception of SST and SST.grd, which showed negative for Krill.
@@
SST.grd 0.0691-0.0691 0.0318-0.0318 $0.0090$
To validate the spatial correlation assumptions underlying each model, we conduct a posterior predictive check (PPC), which assesses model fit by comparing summary statistics of observed data with those of data generated from the posterior predictive distribution. We use Moran’s $I$ statistic (Table 1) as the summary statistic. The PPC procedure is as follows: (i) compute Moran’s $I$ for the observed ranks of each variable; (ii) for each posterior sample $(\bm{R}^{(s)},\phi^{(s)})$, generate replicated latent Gaussian data from (3), i.e., $\bm{z}^{(s)}\sim N_{pn}(\bm{0},\bm{H}(\phi^{(s)})\otimes\bm{R}^{(s)})$, and compute the ranks; (iii) compute Moran’s $I$ for each replicated rank data; (iv) compare the observed Moran’s $I$ with the distribution of the replicated values.
The results reveal a stark contrast between the two models. BGC generates data with virtually no spatial correlation ($I\approx 0$), completely failing to reproduce the strong spatial patterns observed in the data (Table 1). In contrast, spBGC generates spatially correlated data with $I\approx 0.98$. For Depth and SST.grd, the observed Moran’s $I$ falls within the range of replicated values from spBGC, indicating adequate model fit. However, for Krill and Slope, spBGC overestimates spatial correlation, while for SST, it slightly underestimates. This suggests that a single spatial range parameter $\phi$ shared across all variables may not fully capture the varying degrees of spatial dependence, particularly for biological variables (Krill) that exhibit weaker spatial correlation than environmental variables. We also conduct PPC using Matérn 3/2 and Matérn 5/2 correlation functions, with similar results reported in Supplementary Material Section B.
6 Discussion
@@
The proposed method can be applied and extended in several ways. The first is a Gaussian process formulation for the latent variables, from which we can obtain the posterior predictive distribution of the latent variables in arbitrary spatial locations (including non-sampled locations). Then, using the marginal distribution of each outcome estimated in a parametric or non-parametric manner, we can perform spatial predictions in non-sampled locations. Second, although this study focuses on the latent Gaussian process, the proposed method may be conceptually extended to other copula models, such as elliptical copulas and skew elliptical copulas (Smith, 2023). For example, if one is interested in the tail dependence of a latent process, it may be preferable to consider a $t$-copula model. Such heavy-tailed copulas could be particularly useful when dealing with zero-inflated count data, as is common in ecological applications, where extreme co-occurrences may exhibit stronger dependence than what a Gaussian copula can capture. However, posterior computation of additional parameters using such copula models may not be feasible. In this case, an MCMC sampler based on the Metropolis-Hastings algorithm or more efficient computational algorithms may need to be developed. Third, we extend our method to spatiotemporal data. Multivariate data may contain time and location information. For example, the data used in the application example were observed in the Southern Ocean over different periods. In the analysis of the dependence structure of such data, it may be necessary to consider both spatial and temporal correlations. To this end, we consider a hierarchical Bayesian spatiotemporal model based on extended rank likelihood. However, the development of posterior computation for the model is challenging, thus, a subject to address in future work. Finally, although the proposed methodology is for point-referenced data, it is possible to develop a similar method to estimate the copula for multivariate data observed on a graph.